OpenStudy (anonymous):

3s+6 < with a line under it -5(s+2)

OpenStudy (diyadiya):

\[3s+6 \le-5(s+2)\]\[3s+6 \le -5s-10\]\[3s + 5s \le -10-6\]\[8s \le -16\]\[s \le -2\]